OpenStudy (anonymous):

Please help find the answer :) ....√8/9

OpenStudy (lgbasallote):

\[\frac{\sqrt 8}{9}\] what about it?

OpenStudy (anonymous):

I'm supposed to find the value of it and i have the multiple choices for the answer, but I'm confused

OpenStudy (anonymous):

I think the answer is 2sqrt2/3, but I am not sure

OpenStudy (lgbasallote):

first is it \[\frac{\sqrt 8}{9} \; \text{or} \; \sqrt{\frac{8}{9}}\]

OpenStudy (anonymous):

the second one

OpenStudy (lgbasallote):

ahh

OpenStudy (unklerhaukus):

\[\sqrt{\frac{8}{9}}=\frac{\sqrt {8}}{\sqrt{9}} \]

OpenStudy (lgbasallote):

\[\sqrt{\frac 89} \implies \frac{\sqrt 8}{\sqrt 9}\]

OpenStudy (anonymous):

I'm just really confused.... do I multiply them?

OpenStudy (lgbasallote):

actually you're right

OpenStudy (unklerhaukus):

\[8=2\times2\times2=2\times2^2=2^3\] \[9=3\times3=3^2\]

OpenStudy (lgbasallote):

your answer a while ago i mean

OpenStudy (anonymous):

ahaha wait which one?

OpenStudy (unklerhaukus):

\[\frac{\sqrt{8}}{\sqrt{9}}=\frac{\sqrt{2\times2^2}}{\sqrt{3^2}}=\]

OpenStudy (unklerhaukus):

\[=\frac{2\sqrt{2}}{3}.\]

OpenStudy (anonymous):

2x^2sqrt2/3?

OpenStudy (lgbasallote):

x^2? no

OpenStudy (unklerhaukus):

yep

OpenStudy (anonymous):

oh ok yeah my answer was correct

OpenStudy (anonymous):

just saw my mistake thank you

OpenStudy (theviper):

\[\Huge{\color{green}{2\sqrt{2}\over{9}}}\]